# appendices

### v1

@@ Abstract
Figure 1: Percentage of buildings likely damaged within the first two years of the war, aggregated by hromadas. The predictions were thresholded at 0.64, and only buildings larger than 50 m2 were considered.
@@
Nowadays, the analysis of satellite imagery with machine learning offers opportunities to automate parts of the labor-intensive remote monitoring process. Multiple studies have demonstrated the efficiency and effectiveness of this approach for assessing building damage. The most accurate methods currently involve applying deep neural networks to VHR optical data [3, 4, 5, 6], often leveraging large dedicated datasets [7]. These approaches primarily focus on natural disasters, which typically require a one-time analysis of relatively localized, clustered damages. In contrast, armed conflicts can last for months, years, or even decades, resulting in spatially dispersed and gradually increasing destruction. Consequently, they require continuous monitoring over extended periods, so that automated change detection can on the one hand bring larger benefits, but on the other hand is more challenging [8].
@@
Recent work has begun to apply research specifically to the task of assessing building destruction induced by armed conflicts. In a seminal study, Mueller et al. [9] trained a CNN with VHR optical imagery and labels from UNOSAT [10] to assess war-induced building destruction in Syria, leveraging the persistent nature of war damage over time due to the absence of reconstruction efforts amidst ongoing fighting. However, accessing VHR optical imagery for conflict zones is costly, as commercial sources typically do not provide free access to their database in these situations [11], unlike after natural disasters [12, 13]. As such, relying on repeated, large-area coverage with VHR images to monitor armed conflicts at scale is unrealistic for most actors, which recently has prompted studies based on moderate-resolution imagery [14].
Beyond optical images, Synthetic Aperture Radar (SAR) imagery offers a promising alternative. SAR is an active sensor system that illuminates the Earth’s surface with microwave pulses and captures the backscattered signals. It is sensitive to geometrical changes on the ground, such as destroyed buildings. Unlike its optical counterpart, SAR can operate at any time of day (respectively, night) and is largely unaffected by clouds. SAR-based change detection can be divided into amplitude-based methods such as the recent PWTT [15], and phase-based methods such as the popular coherence change detection (CDD) [16, 17]. Due to the sub-meter wavelength of SAR, phase-based methods can detect changes much smaller than the pixel resolution [18]. On the downside, these techniques rely heavily on stable pre-event coherence, which can be problematic with practical, operational satellite SAR systems. Furthermore, creating reliable coherence maps is a complex process that involves a loss of resolution.
@@
In this work, we introduce an easy-to-use, open-source war impact mapping tool based on SAR amplitude data and demonstrate its effectiveness through a comprehensive building damage assessment across Ukraine. Our solution leverages existing, point-wise damage maps from UNOSAT [10] and public, open SAR imagery from the Sentinel-1 mission [19]. Harnessing the parallel computing capabilities of Google Earth Engine (GEE) [20], we train a model to estimate the likelihood of war-related changes from paired time series of Sentinel-1 backscatter. We choose a fixed one-year interval in 2020 as our reference period and iteratively generate damage likelihood maps for 3-month periods ranging from February 2021 up to February 2024. Our model, based on the Random Forest algorithm, allows for straightforward deployment on GEE. In a second processing step, we intersect the per-pixel maps with publicly available building footprints from Overture Maps [21] to generate a damage estimate per building. Our tool produces maps of war damage at very low cost and enables continuous country-scale monitoring. These maps can serve both as a basis for large-scale, aggregated assessments and as guidance to focus manual verification efforts for individual settlements or buildings. We offer two public online dashboards: In the first one, users can dynamically view and explore the pre-computed damage maps described in the following. The second one allows users to run our method themselves for their desired locations and time periods and create custom maps.
The key advantages of our approach are its scalability and ability to transfer to different geographic contexts. Using commercial VHR imagery for regular screening across entire conflict zones is not viable for humanitarian organizations [22], because it would be too resource-intensive, in terms of both direct data cost as well as infrastructure for data download, storage, and compute. By utilizing moderate-resolution satellite images that are free of charge and acquired with short, regular revisits, we enable a more sustainable monitoring approach. The use of SAR imagery has the additional advantage that it can be expected to be transferable to other conflict contexts, including those with different architectural patterns and/or with even more frequent cloud cover. The proposed method is also fast, reproducible, and easy to adapt, even for users who are not experts in remote sensing or image analysis. We believe that our work could be a valuable and accessible resource and could be used to screen large conflict areas beyond our specific application case.
@@ Results
Figure 2: Building damage estimate for Chernihiv (standard confidence threshold 0.64), aggregated over the first two years of conflict. For clarity, we only present the damage heatmap in the main figure, while three zoomed insets show building-level predictions. Red dots represent UNOSAT annotations indicating buildings marked as either destroyed or severely damaged. The VHR satellite layer is displayed solely for visualization, all results are derived from 10m-resolution Sentinel-1 images. Readers are encouraged to explore the maps on the interactive dashboard. Sources: Google Earth/Maxar satellite imagery, Overture Maps building footprints, UNITAR/UNOSAT damage annotations.
Label Precision Recall F1 F0.2 Accuracy AUC
Damaged 69.7 (88.9) 78.7 (40.9) 73.9 (56.0) 70.0 (85.1) 80.7 (77.7) 80.3 (69.1)
Undamaged 87.8 (75.6) 81.8 (97.3) 84.7 (85.1) 87.6 (76.3)
Table 1: Quantitative performance with respect to UNOSAT labels (at previously unseen locations), using a confidence threshold of 0.5. Values in parentheses indicate performance at the optimal threshold of 0.64, maximizing the F0.2 score for the damaged class. The undamaged class is evaluated using the same locations, but acquisition windows before the full-scale invasion. All metrics are given as percentages.
We first evaluate our model’s predictions against additional damage assessments excluded from the training set (see Fig. 4 for the geographic distribution of our test set). The UNOSAT-labeled locations are compared to the corresponding raw pixel predictions of the model. True positive and false negative rates are counted in the 3-month periods from 2022 (after the invasion), and true negative and false positive rates are counted in the periods from 2021 (before the invasion). To guard against potential shifts between the point labels and the satellite tiles, e.g. due to high incidence angles or inaccurate annotation, we compare to the maximum predicted damage probability in a 3$\times$3 pixel window centered at the reference label. We obtain an F1-score of 73.9% for the damaged class and an AUC of 80.3%. As a baseline we also show results for the pixel-wise T-test (PWTT) method [15], using the same evaluation protocol. Our method consistently outperforms that baseline, see Table 2. For details about the comparison see Appendix B.
@@
[…]
$\text{F}_{\beta}=\frac{(1+\beta^{2})\cdot\text{precision}\cdot{}\text{recall}}{\beta^{2}\cdot{}\text{precision}+\text{recall}}$ (1)
In the real-world setting, most pixels are undamaged even in heavily affected locations, we optimize the F0.2-score. With this choice, we find an optimal confidence threshold of 0.64. For details about the threshold optimization see Appendix C. With that value more false positives are suppressed and precision for the damaged class reaches 0.889, at the cost of a lower recall of 0.409. We point out that there is an inherent trade-off between missed detections and false alarms, and the right threshold depends on the application task. Our dashboard allows the user to adjust it with a slider and explore how different thresholds affect the estimates. Table 1 reports results for the settings described above, with a confidence threshold of 0.5 and 0.64.
@@
We use Chernihiv, one of the cities in our test set, to qualitatively illustrate the performance and limitations of our approach. Fig. 2 shows both heatmaps of damage probability and per-building maps after post-processing with building footprints. Inset A highlights one limitation of using building footprints, as the leftmost damage label has been correctly predicted in the heatmap but does not have a corresponding footprint. Inset B illustrates how the model can detect correct patterns of destruction, even for small structures. However, some labels (e.g., center-top) are completely missed, likely because the affected building volume was too small to be visible at Sentinel-1 resolution. Inset C shows both how multiple UNOSAT labels can fall within a single footprint, and how taking the mean prediction over a partially affected building footprint can discard some correct predictions (third point from the bottom). In the main figure, outside the boundaries of UNOSAT analysis, we observe agricultural fields wrongly identified as damaged in the heatmap. This is expected, since our model has not been trained for patterns that occur outside of settlement areas, and illustrates the need to post-process the raw heatmaps with settlement masks or building footprints.
@@
To illustrate the scalability of our approach, we have run damage assessments over the entire country. Our analysis indicates that over 500,000 buildings in Ukraine, or approximately 3.5% of all buildings, have likely sustained damage during the first two years of conflict. This estimate only accounts for buildings larger than 50 m2. Fig. 1 shows the percentage of buildings likely damaged within the first two years of the war, aggregated by hromadas, which is the finest administrative division in Ukraine that covers the entire territory. For the most impacted hromadas with over 10,000 buildings—Bakhmutska, Marinska, and Popasnianska—damage rates reach 47.1%, 32.1%, and 29.0%, respectively. We observe that this pattern of destruction correlates with the course of the war, the eastern side of the country having been a battlefield for months, while the western part has been relatively spared. For a subset of $\approx$1.8M buildings in our dataset, we were able to retrieve meta-data about the building function (such as ”residential” or ”medical”) from the OpenStreetMap catalog. Table 3 summarizes the damaged fractions for different building functions. In this preliminary analysis, we did not find any country-wide patterns that would suggest that certain building types are more prone to war damage than others. Note a more detailed study is needed to confirm or reverse this finding, as biases in the definition and distribution of the available OpenStreetMap building function labels could influence the result.
@@
Residential 70,925 4.5
Industrial 4,207 6.6
Outbuilding 851 2.3
Commercial 1,576 4.7
Education 765 4.0
Agricultural 1,127 6.0
Service 368 4.3
Religious 186 2.3
Medical 152 4.3
Civic 145 5.7
Transportation 133 6.2
Entertainment 11 3.5
Military 7 4.8
@@ Discussion
While so far only tested in Ukraine, we expect our method to transfer well to other geographical contexts, due to the properties of both the underlying Sentinel-1 SAR data and the Random Forest classifier used to analyze it.
@@
Choice of Reference Period: To have a common pre-event reference for both negative and positive samples, we used the year 2020 as a fixed reference period throughout this work. This means, however, that the temporal gap between the pre-event and post-event periods can be rather large; giving rise to mis-classifications due to surface changes other than war damage, e.g., construction. Ideally, a future version of the tool would integrate context knowledge about the conflict to minimize the temporal gap between the pre- and post-event observation periods.
Reference Data: Finally, our supervised learning approach necessitates a sufficient volume of reference data. Despite existing calls for large-scale datasets and open data programs dedicated to armed conflicts [11], such resources remain limited. Data availability varies significantly between different contexts: for some conflicts, e.g. the wars in Ukraine and Gaza, quite a lot of damage data is available; whereas there may be very little information about building damages in conflicts that receive less public attention. It is precisely in these under-reported conflicts that automated analysis could have the biggest impact. By offering an accessible open-access tool based on public satellite data, we hope to take one step toward a global monitoring solution. But to get there it will likely be necessary to collect at least some training data that represent the building patterns and imaging conditions of regions with under-reported conflicts.
@@
But there is inherent uncertainty in predicting damage probabilities from SAR image patterns at 10 m resolution, and users should understand this and align their expectations with what such methods can deliver. Our approach is perhaps best suited to obtain a high-level overview of how damage spreads across time and space. Beyond scientific goals like tracing the development of the conflict, this should often be enough to understand where populations are affected. At the same time, our method cannot confirm damage to an individual building, meaning that it is not directly comparable to assessments based on VHR imagery, ground-based photography, witness reports, etc. Importantly, the automatic large-scale approach and reliable local evidence are complementary: satellite damage maps can identify regions that warrant the effort to gather detailed, reliable evidence.

### v2

@@ Abstract
Figure 1: Percentage of buildings likely damaged within the first two years of the war, aggregated by hromadas. The predictions were thresholded at 0.655, and only buildings larger than 50 m2 were considered.
@@
Nowadays, the analysis of satellite imagery with machine learning offers opportunities to automate parts of the labor-intensive remote monitoring process. Multiple studies have demonstrated the efficiency and effectiveness of this approach for assessing building damage. The most accurate methods currently involve applying deep neural networks to VHR optical data [3, 4, 5, 6, 7, 8], often leveraging large dedicated datasets [9]. These approaches primarily focus on natural disasters, which typically require a one-time analysis of relatively localized, clustered damages. In contrast, armed conflicts can last for months, years, or even decades, resulting in spatially dispersed and gradually increasing destruction. Consequently, they require continuous monitoring over extended periods, so that automated change detection can on the one hand bring larger benefits, but on the other hand, is more challenging  [10].
@@
Recent work has begun to apply research specifically to the task of assessing building destruction induced by armed conflicts. In a seminal study, Mueller et al. [11] trained a CNN with VHR optical imagery and labels from UNOSAT [12] to assess war-induced building destruction in Syria, leveraging the persistent nature of war damage over time due to the absence of reconstruction efforts amidst ongoing fighting. However, accessing VHR optical imagery for conflict zones is costly, as commercial sources typically do not provide free access to their database in these situations [13], unlike after natural disasters [14, 15]. Despite being an active field of research [16, 17], relying on repeated, large-area coverage with VHR images to monitor armed conflicts at scale is not realistic for most actors. Very recently, Hou et al. [18] proposed a deep learning-based approach that can handle both VHR and moderate-resolution optical imagery.
Beyond optical images, Synthetic Aperture Radar (SAR) imagery offers a promising alternative. SAR is an active sensor system that illuminates the Earth’s surface with microwave pulses and captures the backscattered signals. Unlike its optical counterpart, SAR can operate at any time of day (night, respectively) and is largely unaffected by clouds. SAR has a well-established history in damage assessment for both natural and anthropogenic events [19, 20]. Broadly, SAR-based change detection can be divided into incoherent and coherent methods. Incoherent approaches rely only on backscatter amplitude, typically comparing intensity or correlation differences between pre- and post-event images [21, 22, 23, 24, 25]. Coherent methods, on the other hand, use both amplitude and phase information through interferometry (InSAR), generally resulting in more accurate results [26, 27, 28]. Additionally, SAR’s sub-meter wavelength allows coherent techniques to detect changes far smaller than the pixel resolution, particularly when using permanent scatterers, which maintain coherence over long time series [29]. Coherent change detection is a standard approach in the InSAR community to assess building damage after a disaster rapidly [30, 31, 32, 33], c.f. the popular (and patented) Damage Proxy Maps. Recent works have combined machine learning and coherence data to produce more accurate maps, either with handcrafted features [34, 35] or by using recurrent neural networks trained on coherence time series [36, 37]. On the downside, generating high-quality coherence maps is complex and involves a loss of resolution. In this work, we focus on large-scale deployment on cloud-based platforms such as Google Earth Engine (GEE) [38], which do not support the Sentinel-1 SLC product [39] necessary for phase-based methods. We therefore opted for a purely amplitude-based approach.
A few studies have already explored satellite-based damage assessment for the Ukraine War, but most focused on specific areas or events. Aimaiti et al. achieved 58% accuracy on UNOSAT labels in the Kyiv region based on handcrafted features from Sentinel-1 and Sentinel-2 [40]. Huang et al. used optical and coherence SAR time series to identify burnt areas, respectively urban destruction in Mariupol, reaching an overall accuracy of $\approx$60across four classes relative to labels retrieved from social media images [41]. Tavakkoliestahbanati et al. leverage multi-temporal InSAR to monitor deformation from a single event, the Kakhovka dam collapse [42]. Ballinger’s method is arguably the one most directly comparable to ours. It relies on a pixel-wise statistical $t$-test on Sentinel-1 time series, implemented directly in GEE. UNOSAT labels from various locations were then used to calibrate the decision threshold [43]. Lastly, Scher and Van Den Hoek have generated nationwide coherence maps and performed coherent change detection against a reference pre-invasion image. At the time of writing their maps have not yet been made public [44].
@@
In this work, we introduce an easy-to-use, open-source war impact mapping tool based purely on SAR amplitude data and demonstrate its effectiveness through a comprehensive building damage assessment across Ukraine. Our solution leverages existing, point-wise damage maps from UNOSAT [12] and public, open SAR imagery from the Sentinel-1 mission [39]. Harnessing the parallel computing capabilities of GEE [38], we train a model to estimate the likelihood of war-related changes from paired time series of Sentinel-1 backscatter. We choose a fixed one-year interval in 2020 as our reference period and iteratively generate damage likelihood maps for 3-month periods ranging from February 2021 up to February 2024. Our model, based on the Random Forest algorithm, allows for straightforward deployment on GEE and outputs damage maps with 10$\times$10 m grid spacing, the original ground sampling distance (GSD) of the Sentinel-1 GRD product. In a second processing step, we intersect the per-pixel maps with publicly available building footprints from Overture Maps [45] to generate a damage estimate per building. Our tool produces maps of war damage at a very low cost and enables continuous country-scale monitoring. These maps can serve both as a basis for large-scale, aggregated assessments and as guidance to focus manual verification efforts for individual settlements or buildings. We offer two public online dashboards: In the first one, users can dynamically view and explore the pre-computed damage maps described in the following. The second one allows users to run our method themselves for their desired locations and time periods and create custom maps.
The key advantages of our approach are its scalability and ability to transfer to different geographic contexts. Using commercial VHR imagery for regular screening across entire conflict zones is not viable for humanitarian organizations [46], because it would be too resource-intensive, in terms of both direct data cost as well as infrastructure for data download, storage, and compute. By utilizing moderate-resolution satellite images that are free of charge and acquired with short, regular revisits, we enable a more sustainable monitoring approach. The use of SAR imagery has the additional advantage that the same method can be expected to work well in other conflict contexts, including those with different architectural patterns and/or with even more frequent cloud cover, as it is purely geometrical and does not rely on any location-specific a-priori assumptions. The proposed method is also fast, reproducible, and easy to adapt, even for users who are not experts in remote sensing or image analysis. We believe that our work could be a valuable and accessible resource and could be used to screen large conflict areas beyond our specific application case.
@@ Results
Figure 2: Building damage estimate for Chernihiv (standard confidence threshold 0.655), aggregated over the first two years of conflict. For clarity, we only present the damage heatmap in the main figure, while three zoomed insets show building-level predictions. Red dots represent UNOSAT annotations indicating buildings marked as either destroyed or severely damaged. The VHR satellite layer is displayed solely for visualization, all results are derived from 10 m-resolution Sentinel-1 images. Readers are encouraged to explore the maps on the interactive dashboard. Sources: Google Earth/Maxar satellite imagery, Overture Maps building footprints, UNITAR/UNOSAT damage annotations.
Label Precision Recall F1 Accuracy AUC
Damaged 67.1 (90.3) 84.6 (46.5) 74.9 (61.4) 80.3 (79.7) 81.3 (71.9)
Undamaged 90.5 (77.4) 78.0 (97.3) 83.8 (86.2)
Table 1: Quantitative detection performance relative to UNOSAT labels (at previously unseen locations), using a confidence threshold of 0.5. Values in parentheses indicate performance at the threshold of 0.655, reaching a target precision of 90%. The undamaged class is evaluated using the same locations, but acquisition windows before the full-scale invasion. All metrics are given as percentages.
We first evaluate our model’s predictions against additional damage assessments excluded from the training set (see Fig. 4 for the geographic distribution of our test set). The UNOSAT-labeled locations are compared to the corresponding raw pixel predictions of the model. True positive and false negative rates are counted in the 3-month periods from 2022 (after the invasion), and true negative and false positive rates are counted in the periods from 2021 (before the invasion). To guard against potential shifts between the point labels and the satellite tiles, e.g. due to high incidence angles or inaccurate annotation, we compare to the maximum predicted damage probability in a 3$\times$3 pixel window centered at the reference label. We report our results in Table 1. We obtain an F1-score of 74.9% for the damaged class and an AUC of 81.3%. Using the same evaluation protocol, we compare our result to the pixel-wise $t$-test (PWTT) method [43]. To our knowledge, this is the only method directly comparable to ours, since others either require coherence maps or assume that the precise date of the destruction event is known, which is not the case for our UNOSAT labels. Our method consistently outperforms that baseline, see Table 2. For details about the comparison see Appendix B.
@@
The evaluation of held-out validation data also helps establish an appropriate threshold for distinguishing between damaged and undamaged pixels. Indeed, while the model has been trained on a balanced dataset, the actual distribution of damaged versus undamaged pixels is highly skewed, and the prevalence of undamaged pixels in the test set means that thresholding at damage probability 0.5 will cause a significant number of false alarms. To address this, we set a target precision of 90%, aligning with the reality of operational deployment, where it is crucial to keep the false positive rate at a manageable level. With this choice, we find an optimal threshold of 0.655. For details about the threshold optimization see Appendix C. With that value, the recall reaches 46.5%. We point out that there is an inherent trade-off between missed detections and false alarms, and the right threshold depends on the application task. Our dashboard allows the user to adjust it with a slider and explore how different thresholds affect the estimates. Results with the confidence threshold of 0.655 are also reported in Table 1.
@@
We use Chernihiv, one of the cities in our test set, to qualitatively illustrate the performance and limitations of our approach. Fig. 2 shows both heatmaps of damage probability and per-building maps after post-processing with building footprints. Inset A highlights one limitation of using building footprints, as the leftmost damage label has been correctly predicted in the heatmap but does not have a corresponding footprint. Inset B illustrates how the model can detect correct patterns of destruction, even for small structures. However, some labels (e.g., center-top) are completely missed, likely because the affected building volume was too small to be visible at Sentinel-1 resolution. Inset C shows both how multiple UNOSAT labels can fall within a single footprint. In the main figure, outside the boundaries of UNOSAT analysis, we observe agricultural fields wrongly identified as damaged in the heatmap. This is expected, since our model has not been trained for patterns that occur outside of settlement areas, and illustrates the need to post-process the raw heatmaps with settlement masks or building footprints.
@@
To illustrate the scalability of our approach, we have run damage assessments over the entire country. Our analysis indicates that over 400,000 buildings in Ukraine, or approximately 2.7% of all buildings, have likely sustained damage during the first two years of conflict. This estimate only accounts for buildings larger than 50 m2. Fig. 1 shows the percentage of buildings likely damaged within the first two years of the war, aggregated by hromadas, which is the finest administrative division in Ukraine that covers the entire territory. For the most impacted hromadas with over 10,000 buildings—Bakhmutska, Marinska, and Popasnianska—damage rates reach 42.6%, 29.0%, and 25.0%, respectively. We observe that this pattern of destruction correlates with the course of the war, the eastern side of the country having been a battlefield for months, while the western part has been relatively spared. For a subset of $\approx$1.8M buildings in our dataset, we were able to retrieve meta-data about the building function (such as ”residential” or ”medical”) from the OpenStreetMap catalog. Table 3 summarizes the damaged fractions for different building functions. In this preliminary analysis, we did not find any country-wide patterns that would suggest that certain building types are more prone to war damage than others. Note a more detailed study is needed to confirm or reverse this finding, as biases in the definition and distribution of the available OpenStreetMap building function labels could influence the result.
@@
Residential 57,089 3.6
Industrial 3,177 5.0
Outbuilding 599 1.6
Commercial 1,309 3.9
Education 658 3.5
Agricultural 775 4.1
Service 284 3.6
Religious 177 2.1
Medical 126 3.6
Civic 117 4.6
Transportation 106 4.9
Entertainment 8 2.6
Military 6 4.1
@@ Discussion
While our method has so far been tested only in Ukraine, we plan to extend its application to other regions. Given the characteristics of both the Sentinel-1 SAR data and the Random Forest classifier, we anticipate that the methodology will adapt well to new areas, as long as we have adequate data for fine-tuning. However, the final performance will ultimately be influenced by factors such as vegetation cover and building density.
@@
Choice of Reference Period: To have a common pre-event reference for both negative and positive samples, we used the year 2020 as a fixed reference period throughout this work. This means, however, that the temporal gap between the pre-event and post-event periods can be rather large; giving rise to misclassifications due to surface changes other than war damage, e.g., construction. Ideally, a future version of the tool would integrate context knowledge about the conflict to minimize the temporal gap between the pre- and post-event observation periods.
Using Sentinel-1 data from GEE: Our study focuses on large-scale assessment and takes full advantage of GEE’s computational capabilities. However, it is important to note that while the Sentinel-1 data processed by GEE are orthorectified, they are not radiometrically terrain flattened. This limitation is particularly relevant in areas with significant slope. While most cities in Ukraine are on fairly flat terrain, our method may not perform as well in towns and cities located in more mountainous areas. Additionally, as previously mentioned, GEE can at present not ingest complex data, which prevents the use of phase coherence for change detection. In future research, we plan to explore alternative approaches, so as to fully exploit the potential of SAR data.
Reference Data: Our supervised learning approach necessitates a sufficient volume of reference data. Despite existing calls for large-scale datasets and open data programs dedicated to armed conflicts [13], such resources remain limited. Data availability varies significantly between different contexts: for some conflicts, e.g. the wars in Ukraine and Gaza, quite a lot of damage data is available; whereas there may be very little information about building damages in conflicts that receive less public attention. It is precisely in these under-reported conflicts that automated analysis could have the biggest impact. By offering an accessible open-access tool based on public satellite data, we hope to take one step toward a global monitoring solution. But to get there it will likely be necessary to collect at least some training data that represent the building patterns and imaging conditions of regions with under-reported conflicts.
@@
But there is inherent uncertainty in predicting damage probabilities from SAR image patterns at 10 m GSD, and users should understand this and align their expectations with what such methods can deliver. Our approach is perhaps best suited to obtain a high-level overview of how damage spreads across time and space. Beyond scientific goals like tracing the development of the conflict, this should often be enough to understand where populations are affected. At the same time, our method cannot confirm damage to an individual building, meaning that it is not directly comparable to assessments based on VHR imagery, ground-based photography, witness reports, etc. Importantly, the automatic large-scale approach and reliable local evidence are complementary: satellite damage maps can identify regions that warrant the effort to gather detailed, reliable evidence.